OpenStudy (anonymous):

Ben drives 405 miles on 15 gallons. How many miles does Ben drive on one gallon?

OpenStudy (anonymous):

the answers are 4.05 miles per gallon 27 miles per gallon 1 over 27. miles per gallon 405 miles per 15 gallon

OpenStudy (alexandervonhumboldt2):

x/1 gallon=405/15 x=405/15=solve

OpenStudy (anonymous):

thanks